# Sparse Autoencoders Can Interpret Randomly Initialized Transformers

### v1

@@ Abstract
Sparse autoencoders (SAEs) are an increasingly popular technique for interpreting the internal representations of transformers. In this paper, we apply SAEs to ‘interpret’ random transformers, i.e., transformers where the parameters are sampled IID from a Gaussian rather than trained on text data. We find that random and trained transformers produce similarly interpretable SAE latents, and we confirm this finding quantitatively using an open-source auto-interpretability pipeline. Further, we find that SAE quality metrics are broadly similar for random and trained transformers. We find that these results hold across model sizes and layers. We discuss a number of number interesting questions that this work raises for the use of SAEs and auto-interpretability in the context of mechanistic interpretability.
@@ 1 Introduction
Sparse autoencoders (SAEs) have become a popular tool in mechanistic interpretability research, with the aim of disentangling the internal representations of neural networks by learning sparse, interpretable features from network activations (Elhage et al., 2022; Sharkey et al., 2022; Cunningham et al., 2023; Bricken et al., 2023). An autoencoder with a high-dimensional hidden layer is trained to reconstruct activations while enforcing sparsity (Gao et al., 2024; Templeton et al., 2024; Lieberum et al., 2024), with the aim of discovering the underlying concepts or ‘features’ learned by the network (Park et al., 2023; Wattenberg & Viégas, 2024).
An interpretability technique that aims to understand the computations a model has learned to perform during training should distinguish between trained and untrained models or models whose parameters have been randomized. We show that SAEs do not always achieve this, suggesting they provide limited insights into the computational structure learned by the underlying model. In particular, we found that the latents learned by SAEs trained on transformers whose parameters are randomized have similar auto-interpretability scores (Bills et al., 2023; Paulo et al., 2024) to the latents learned by SAEs trained on the activations of the corresponding unmodified transformer. Additionally, we found that randomized transformers are much closer to trained models than to our control in terms of standard SAE evaluations.
This result raises important questions about what we can glean from applying SAEs to interpret language models. While SAEs extract interpretable features (e.g., Cunningham et al., 2023; Bricken et al., 2023; Rajamanoharan et al., 2024a), our findings suggest that many features are likely to reflect statistical properties of the training data (Dooms & Wilhelm, 2024) and underlying model architecture rather than learned computations. This challenges the widespread interest in understanding the mechanisms of trained models with SAEs (Sharkey et al., 2025), highlighting the need to develop techniques that better capture computations.
@@
These results have important implications for mechanistic interpretability research. In particular, we suggest that more rigorous methods to distinguish between artifacts and genuinely learned computations are needed, and that interpretability techniques should be carefully validated against appropriate null models.
Finally, we speculate about why these patterns might emerge. At a high level, there are two hypotheses:
  1. 1.
    The input data is already superposed, and randomly initialized NNs largely preserve this superposition.
  2. 2.
    Randomly initialized NNs amplify or even introduce superposed structure to the input data (e.g., given dense input generated IID from a Gaussian).
We present toy models to demonstrate the plausibility of these hypotheses in Section 4 but defer conclusions as to the mechanism responsible to future work.
@@ 2 Related Work
Under a different name, ‘superposition’ in visual data is one of the foundational observations of computational neuroscience. Olshausen & Field (1996, 1997) showed that the receptive fields of simple cells in the mammalian visual cortex can be explained as a result of sparse coding, i.e., representing a relatively large number of signals (sensory information) by simultaneously activating a relatively small number of elements (neurons). Coding theory offers a perspective on efforts to extract the ‘underlying signals’ responsible for neural network activations (Marshall & Kirchner, 2024).
Sparse dictionary learning (SDL) approximates a set of input vectors by linear combinations of a relatively small number of learned basis vectors. The learned basis is usually overcomplete: it has a greater dimension than the inputs. SDL algorithms include Independent Component Analysis (ICA), which finds a linear representation of the data such that the components are maximally statistically independent (Bell & Sejnowski, 1995; Hyvärinen & Oja, 2000). Sparse autoencoders (SAEs) are a simple neural network approach (Lee et al., 2006; Ng, 2011; Makhzani & Frey, 2014). Typically, an autoencoder with a single hidden layer that is many times larger than the input activation vectors is trained with an objective that imposes or incentivizes sparsity in its hidden layer activations to try to find this structure.
@@
Recently, it has become common to understand ‘features’ or concepts in language models as low-dimensional subspaces of internal model activations (Park et al., 2023; Wattenberg & Viégas, 2024; Engels et al., 2024). If such sparse or ‘superposed’ structure exists, we expect to be able to ‘intervene on’ or ‘steer’ the activations, i.e., to modify or replace them to express different concepts and so influence model behavior (Meng et al., 2022; Zhang & Nanda, 2023; Heimersheim & Nanda, 2024; Makelov, 2024; O’Brien et al., 2024).
SAEs are a popular approach to discovering these features, where we typically train a single autoencoder to reconstruct the activations of a single neural network layer, e.g., the transformer residual stream (Sharkey et al., 2022; Cunningham et al., 2023; Bricken et al., 2023). A wide variety of SAE architectures have been suggested, which commonly vary the activation function applied after the linear encoder (Makhzani & Frey, 2014; Gao et al., 2024; Rajamanoharan et al., 2024b; Lieberum et al., 2024). SAEs have also been trained with different objectives (Braun et al., 2024) and applied to multiple layers simultaneously (Yun et al., 2021; Lawson et al., 2024; Lindsey et al., 2024).
Besides reconstruction errors and preservation of the underlying model’s performance, SAEs have been evaluated according to whether they capture specific concepts (Gurnee et al., 2023; Gao et al., 2024) or factual knowledge (Huang et al., 2024; Chaudhary & Geiger, 2024), and whether these can be used to ‘unlearn’ concepts (Karvonen et al., 2024b).
@@
Sparse autoencoders often learn tens of thousands of latents, which are infeasible to describe by hand. Yun et al. (2021) find the tokens that maximally activate a dictionary element from a text dataset and manually inspect activation patterns. Instead, researchers typically collect latent activation patterns over a text dataset and prompt a large language model to explain them (e.g. Bills et al., 2023; Foote et al., 2023). These methods have been widely adopted (e.g. Cunningham et al., 2023; Bricken et al., 2023; Gao et al., 2024; Templeton et al., 2024; Lieberum et al., 2024).
@@
Lecomte et al. (2024) point out that neurons may become polysemantic incidentally. Polysemanticity is the notion that a single neuron (basis dimension) of a network layer may represent multiple interpretable concepts (Elhage et al., 2022; Scherlis et al., 2023); unsurprisingly, individual neurons in a randomly initialized network may be polysemantic (Lecomte et al., 2024). By contrast, our work studies superposition (Elhage et al., 2022; Chan, 2024), which pertains to the representations learned across a whole network layer as opposed to any individual neuron. In particular, superposition allows a network layer as a whole to represent a larger number of (sparse) features than the layer has (dense) neurons by sparse coding (only a few concepts are active at a time, i.e., a given token position).
@@
Recently, Zhong & Andreas (2024) demonstrated some surprising capabilities of transformers where only the embeddings are trained and no other parameters. These results demonstrate that the behavior of a randomly initialized transformer can be shaped to a surprising extent by training only a few parameters. However, this is very different from our setting, not only in that we consider SAEs, but also in that our ‘Step-0’ and ‘Re-randomized incl. embeddings’ variants randomize all the parameters, including the embeddings. While the ‘Re-randomized excl. embeddings’ variant does use pre-trained embeddings, we do not train those embeddings with fixed, randomized weights but simply freeze the pre-trained embeddings and randomize the other weights (Section 3).
@@ 3 Results
We trained per-layer SAEs on the residual stream activation vectors of transformer language models from the Pythia suite with between 70 million and 7 billion parameters (Biderman et al., 2023). We compared SAEs trained on different variants of the underlying transformers:
@@
Control: The original, trained model, except where the input token embeddings are replaced at inference time by sampling IID standard Gaussian noise for each token, such that a given token does not have a consistent embedding vector. We expect auto-interpretability to perform at the level of chance for this variant.
@@
We train on 100 million tokens from the RedPajama dataset (Weber et al., 2024; TogetherComputer, 2023) with an activation buffer size of 10 million tokens. For models with fewer than 410 million parameters, we trained an SAE at every layer; for Pythia-1B, at every second layer; and for Pythia-6.9B, at every fourth layer.
@@
Unless otherwise stated, we trained $k$-sparse autoencoders, a.k.a. TopK SAEs, with an expansion factor of $R=64$ and $k=32$. We confirm that our results are robust with respect to these hyperparameters by training SAEs on Pythia-160M with expansion factors equal to powers of 2 between 16 and 128 and $k$ values of 16 and 32 (Figure 11).
Our training implementation is based on Belrose (2024); our evaluation implementations are based on EleutherAI and Karvonen et al. (2024a).
3.1 Auto-interpretability
Feature explanations, which identify some concept, can be used as input to a classifier that predicts the presence of that concept within a piece of text. We can evaluate such a classifier using traditional metrics like the area under the receiver operating characteristic curve (AUROC). There is a choice of classification tasks for evaluating feature explanations, e.g., fuzzing and detection scoring (Paulo et al., 2024).
@@ 3.1 Auto-interpretability
We use the ‘fuzzing’ scoring method, where both positive and negative examples of tokens (i.e., with non-zero and zero activation values, respectively) for a given latent are delimited with special characters, and a language model is prompted to identify which examples have been correctly delimited for the latent given its explanation. By contrast, simulation scoring is based on the correlation between simulated and observed activations (Bills et al., 2023).
@@
For each trained SAE (i.e., underlying model, variant, and layer), we randomly sampled 100 features to obtain auto-interpretability scores. The implementation is based on Paulo et al. (2024). We use the Meta-Llama-3.1-70B-Instruct-AWQ-INT4 model to generate explanations and make predictions, larger than the 8B models used by Choi et al. (2024) and open-source unlike Bills et al. (2023).
@@
We found that the auto-interpretability scores were far more similar between the trained and randomized models than with the control (Figures 1, 2, and 3). The similarity between the ROC scores for trained and randomized transformers demonstrates that ‘fuzzing’ auto-interpretability, applied to SAE latent explanations, fails to meaningfully distinguish between these underlying models.
Figure 2: ROC curves for ‘fuzzing’ auto-interpretability for Pythia-6.9B over 100 SAE latents. Similarly to Figure 1, these results demonstrate the similarity in performance between the SAE variants variants. Notably, the trained variant appears to degrade in performance for the later model layers.
3.2 Evaluation
We went on to consider standard SAE quality metrics, along with the auto-interpretability AUROC for Pythia models with between 70M and 6.9B parameters. As above, we broadly found that the randomly initialized and re-randomized models (Figure 3 blue, green, orange lines) were more similar to the trained model (Figure 3 gray lines) than to our control (Figure 3 black lines).
Notably, the cosine similarity and explained variance are often far lower for the random control than the other models, and its reconstruction errors tend to increase across layers while the remaining variants decrease. For the random control, this can perhaps be explained by the fact that a Gaussian is the highest entropy distribution with fixed mean and variance (Jaynes, 2003); we speculate that Gaussian vectors are the ‘least structured’, in some sense, and thus hardest for SAEs to reconstruct. As Gaussian-distributed activations are propagated through successive layers, we would expect the activations to become less Gaussian and perhaps more ‘sparse’, i.e., easier to reconstruct (Section 4).
Interestingly, the randomized variants (blue and orange lines) are more similar to the trained model than the variant at initialization (green line). This is especially evident if we look at the $L^{1}$ norm values in larger models. We speculate that this pattern arises because parameter norms may differ greatly between a trained model and its state at initialization. In contrast, our randomization procedure was specifically designed to preserve parameter norms with respect to the trained model. The scale of parameters at different layers may be important, e.g., to control the growth of activations as they progress through the residual stream (Liu et al., 2020). In the AUROC plots, we find that for all but the control variant, AUROC increases with model size. We speculate that features become more specific as the size of the SAEs increases: in smaller SAEs, each latent has to explain more of the input, making classification tasks easier.
Figure 3 (fifth row) shows the cross-entropy (CE) loss score, a.k.a. the loss recovered, broken down by layer. This is the increase in the CE loss when the original model activations are replaced by their SAE reconstructions, divided by the increase in the CE loss when the activations are replaced by zeros (‘ablated’). The results show that the ‘trained’ variant SAEs perform similarly to others from the literature (e.g., Kissane et al., 2024; Rajamanoharan et al., 2024a; Mudide et al., 2024). Importantly, the CE loss score only makes sense for the trained variant: for any of the randomized variants, the CE loss is very poor regardless of whether the original or reconstructed activations are used.
@@ 3.2 Evaluation
3.3 Latent explanation complexity
@@
While randomized and trained transformers achieve similar auto-interpretability scores and evaluation metrics, we suspected that SAEs applied to these very different settings might capture qualitatively different features. In particular, we suspected that SAEs trained on the randomized variants could learn simple features depending primarily on characteristics of the input text (Dooms & Wilhelm, 2024) but not more complex, abstract features, like those captured with trained transformers (e.g. Templeton et al., 2024). Given the difficulty of qualitatively evaluating features at scale, we attempt a quantitative evaluation in terms of the distribution of latent activations over token IDs. We provide a random sample of features and the corresponding maximally activating dataset examples for each SAE variant of Pythia-6.9B in Appendix D.
@@
Anecdotally, we have observed that a significant proportion of SAE latents have non-zero activations only on a single token or a small number of distinct tokens within a text dataset (e.g., Lin & Bloom, 2024; Dooms & Wilhelm, 2024). Hence, a simple measure of the complexity of an SAE latent given a set of maximally activating examples is the degree to which the latent activates on a single token ID or multiple distinct token IDs. Specifically, we quantify the number of token IDs in terms of the entropy of the observed distribution of latent activations over tokens (vocabulary items): the greater the entropy, the more ‘spread out’ the latent activations over the vocabulary, and the less token-specific the latent. We take this distribution to be the total latent activation per token across the set of maximally activating examples used to generate explanations for auto-interpretability.
@@
We include the entropy of the observed distributions of latent activations over token IDs in the last row of Figure 3. The negative control variant displays a consistently high entropy, which is to be expected given that the embedding for a given token ID is sampled IID from a Gaussian on each occurrence of the token, i.e., a token does not have a consistent embedding vector (Section 3). For the trained variant, the entropy increases across layers, i.e., the further into the model, the less likely the maximally activating examples for each latent contain activations concentrated on a single token. This is also expected: at later layers, we expect more abstract features that are less similar to token embeddings. Finally, the entropy for randomized models tends to be lower than either the trained or control variants, indicating that latents are activated specifically at a single token ID or a small number of IDs.
@@
In combination with the preceding results, this suggests that standard SAE quality and auto-interpretability metrics are missing an important aspect of SAE features: their “abstractness”. While the token distribution entropy is not a direct measure of “abstractness”, it suggests that the randomized variants, viewed in the context of their similar auto-interpretability scores to the trained variant, remain able to learn simple, single-token features. However, unlike the trained variant, the features of the randomized variants do not become more complex as the layer index increases.
4 A toy model of superposition

### v2

# Automated Interpretability Metrics Do Not Distinguish Trained and Random Transformers

@@ Abstract
Sparse autoencoders (SAEs) are widely used to extract sparse, interpretable latents from transformer activations. We test whether commonly used SAE quality metrics and automatic explanation pipelines can distinguish trained transformers from randomly initialized ones (e.g., where parameters are sampled i.i.d. from a Gaussian). Over a wide range of Pythia model sizes and multiple randomization schemes, we find that, in many settings, SAEs trained on randomly initialized transformers produce auto-interpretability scores and reconstruction metrics that are similar to those from trained models. These results show that high aggregate auto-interpretability scores do not, by themselves, guarantee that learned, computationally relevant features have been recovered. We therefore recommend treating common SAE metrics as useful but insufficient proxies for mechanistic interpretability and argue for routine randomized baselines and targeted measures of feature ‘abstractness.’
@@ 1 Introduction
Sparse autoencoders (SAEs) are a popular tool in mechanistic interpretability research, with the aim of disentangling the internal representations of neural networks by learning sparse, interpretable features from network activations (Elhage et al., 2022; Sharkey et al., 2022; Cunningham et al., 2023; Bricken et al., 2023). An autoencoder with a high-dimensional hidden layer is trained to reconstruct activations while enforcing sparsity (Gao et al., 2024; Templeton et al., 2024; Lieberum et al., 2024), with the aim of discovering the underlying concepts or ‘features’ learned by the network (Park et al., 2023; Wattenberg and Viégas, 2024). Developing better SAEs relies on quantitative evaluation metrics like auto-interpretability scores that measure agreement between generated explanations and activation patterns (Bills et al., 2023; Paulo et al., 2024; Karvonen et al., 2024a).
For an interpretability method to be considered robust, its evaluation metrics should distinguish features learned through training from artifacts arising from the data or model architecture. A key sanity check is therefore to compare the method’s output on a trained model against a strong null model, such as one with randomly initialized weights (Adebayo et al., 2020). We apply this sanity check to SAEs and find that several common quantitative metrics do not always clearly distinguish between the trained and randomized settings. In particular, we found that SAEs trained on transformers with random parameters can yield latents with auto-interpretability scores (Bills et al., 2023; Paulo et al., 2024) that are surprisingly similar to those from a fully trained model.
This result raises important questions about what we can glean from applying these metrics of SAE quality. High auto-interpretability scores alone do not guarantee that an SAE has identified complex, learned computations. Instead, such scores may sometimes reflect simpler statistical properties of the training data (Dooms and Wilhelm, 2024) or architectural inductive biases that are present even without training. Indeed, one could argue that a randomly initialized network still performs a basic form of computation, such as preserving or amplifying the sparse structure of its inputs (Section 4). From this perspective, SAEs might faithfully interpret this simple, inherent computation.
@@
While some SAE features from trained models clearly arise from learned computation, the commonly used aggregate metrics are often insufficient for determining whether a given SAE has learned these more complex features. These results have important implications for mechanistic interpretability research. In particular, we suggest that more rigorous methods to distinguish between artifacts and genuinely learned computations are needed, and that interpretability techniques should be carefully validated against appropriate null models.
Finally, we speculate about why these patterns might emerge. At a high level, there are two hypotheses: (1) the input data already exhibits superposition, and randomly initialized neural networks largely preserve this superposition; and (2) randomly initialized neural networks amplify or even introduce superposed structure to the input data (e.g., given dense input generated i.i.d. from a Gaussian). We present toy models to demonstrate the plausibility of these hypotheses in Section 4 but defer conclusions as to the mechanism responsible to future work.
@@ 2 Related Work
Under a different name, ‘superposition’ in visual data is one of the foundational observations of computational neuroscience. Olshausen and Field (1996; 1997) showed that the receptive fields of simple cells in the mammalian visual cortex can be explained as a result of sparse coding, i.e., representing a relatively large number of signals (sensory information) by simultaneously activating a small number of elements (neurons). Coding theory offers a perspective on efforts to extract the ‘underlying signals’ responsible for neural network activations (Marshall and Kirchner, 2024).
Sparse dictionary learning (SDL) approximates a set of input vectors by linear combinations of a relatively small number of learned basis vectors. The learned basis is usually overcomplete: it has a greater dimension than the inputs. SDL algorithms include Independent Component Analysis (ICA), which finds a linear representation of the data such that the components are maximally statistically independent (Bell and Sejnowski, 1995; Hyvärinen and Oja, 2000). Sparse autoencoders (SAEs) are a simple neural network approach (Lee et al., 2006; Ng, 2011; Makhzani and Frey, 2014). Typically, an autoencoder with a single hidden layer that is many times larger than the input activation vectors is trained with an objective that imposes or incentivizes sparsity in its hidden layer activations to try to find this structure. A latent is a single neuron (dimension) in the autoencoder’s hidden layer.
@@
Recently, it has become common to understand ‘features’ or concepts in language models as low-dimensional subspaces of internal model activations (Park et al., 2023; Wattenberg and Viégas, 2024; Engels et al., 2024). If such sparse or ‘superposed’ structure exists, we expect to be able to ‘intervene on’ or ‘steer’ the activations, i.e., to modify or replace them to express different concepts and so influence model behavior (Meng et al., 2022; Zhang and Nanda, 2023; Heimersheim and Nanda, 2024; Makelov, 2024; O’Brien et al., 2024).
SAEs are a popular approach for discovering features, where one typically trains a single autoencoder to reconstruct the activations of a single neural network layer, e.g., the transformer residual stream (Sharkey et al., 2022; Cunningham et al., 2023; Bricken et al., 2023). Many SAE architectures have been suggested, which commonly vary the activation function applied after the linear encoder (Makhzani and Frey, 2014; Gao et al., 2024; Rajamanoharan et al., 2024b; Lieberum et al., 2024). SAEs have also been trained with different objectives (Braun et al., 2024; Farnik et al., 2025) and applied to multiple layers simultaneously (Yun et al., 2021; Lawson et al., 2024; Lindsey et al., 2024).
Besides reconstruction errors and preservation of the underlying model’s performance, SAEs have been evaluated according to whether they capture specific concepts (Gurnee et al., 2023; Gao et al., 2024) or factual knowledge (Huang et al., 2024; Chaudhary and Geiger, 2024), and whether these can be used to ‘unlearn’ concepts (Karvonen et al., 2024b).
@@
SAEs often learn tens of thousands of latents, which are infeasible to describe by hand. Yun et al. (2021) find the tokens that maximally activate a dictionary element from a text dataset and manually inspect activation patterns. Instead, researchers typically collect latent activation patterns over a text dataset and prompt a large language model to explain them (e.g. Bills et al., 2023; Foote et al., 2023). These methods have been widely adopted (e.g. Cunningham et al., 2023; Bricken et al., 2023; Gao et al., 2024; Templeton et al., 2024; Lieberum et al., 2024).
@@
Lecomte et al. (2024) noted that neurons may become polysemantic incidentally. A polysemantic neuron (basis dimension) of a network layer represents multiple interpretable concepts (Elhage et al., 2022; Scherlis et al., 2023); unsurprisingly, individual neurons in a randomly initialized network may be polysemantic. By contrast, our work studies superposition (Elhage et al., 2022; Chan, 2024), which pertains to the representations learned across a whole network layer as opposed to any individual neuron. In particular, superposition allows a network layer as a whole to represent a larger number of (sparse) features than the layer has (dense) neurons by sparse coding (only a few concepts are active at a time, i.e., a given token position).
@@
Zhong and Andreas (2024) showed that transformers learn surprising algorithmic capabilities when only the embeddings are trained and no other parameters. These results demonstrate that the behavior of a randomly initialized transformer can be shaped to a surprising extent by training only a few parameters. However, our setting is very different: besides considering SAEs, we randomize all the parameters, including the embeddings, in our ‘Step-0’ and ‘Re-randomized incl. embeddings’ variants. Our ‘Re-randomized excl. embeddings’ variant uses pre-trained embeddings, but we do not train those embeddings with fixed, randomized weights. Instead, we freeze the pre-trained embeddings and randomize the other weights (Section 3).
Random transformers for board games
Karvonen et al. (2024c) found that SAEs were considerably better at extracting meaningful structure from chess games using pre-trained transformers, as opposed to those with random weights. However, the data from board games is wildly different from language data. In particular, there is reason to expect that language is sparse (e.g., a particular concept such as ‘serendipitous’ appears only rarely), and that this sparse structure is ‘aligned’ with conceptual meaning. In contrast, in board games, this is not necessarily true: a useful concept such as a knight fork does not necessarily turn up sparsely in board games.
Random one-layer transformers
Bricken et al. (2023) found that auto-interpretability scores discriminated effectively between random and trained one-layer transformers. Similarly, we found that auto-interpretability scores for randomized models were relatively low for smaller models (e.g., Pythia-70m) but that the gap was narrowed for larger models (e.g., Pythia-6.9b).
@@ 3 Results
Figure 1: ‘Fuzzing’ ROC curve vs. layer for Pythia-6.9b (100 latents sampled per SAE). The trained model (gray line) and randomized variants (colored) overlap, whereas the control (black) is near chance (dotted). This suggests aggregate AUROC alone is insufficient to attribute latents to learned computation. See Figure 2 for other metrics/model sizes and Appendix E for multiple random seeds.
We trained per-layer SAEs on the residual stream activation vectors of transformer language models from the Pythia suite, with between 70M and 7B parameters (Biderman et al., 2023). We compared SAEs trained on different variants of the underlying transformers:
@@
  • Control: The original, trained model, except where the input token embeddings are replaced at inference time by sampling i.i.d. standard Gaussian noise for each token, such that a given token does not have a consistent embedding vector. For this variant, we expect auto-interpretability to perform at the level of chance.
@@
For our primary experiments, we trained SAEs on 100M tokens from the RedPajama dataset (Weber et al., 2024) using an activation buffer size of 10M tokens (see Appendix C for a subset of experiments that demonstrate similar results with SAEs trained on one billion tokens). For models with fewer than 410M parameters, we trained an SAE at every layer; for Pythia-1b, we trained SAEs at every second layer; and for Pythia-6.9b, we trained SAEs at every fourth layer.
@@
Unless otherwise stated, we trained $k$-sparse autoencoders (also known as TopK SAEs; Makhzani and Frey 2014; Gao et al. 2024), with an expansion factor of $R=64$ and sparsity $k=32$. We confirm that our results are robust with respect to these hyperparameters by training SAEs on Pythia-160m with expansion factors equal to powers of 2 between 16 and 128, and sparsities of 16 and 32 (Figure 18). The training implementation is based on Belrose et al. (2025); our evaluations are based on Caden et al. (2025) and Karvonen et al. (2024a).
Auto-interpretability
@@
Feature explanations that identify a concept can be input to a classifier that predicts whether the concept appears in the text inputs. Such a classifier may be evaluted by traditional metrics, like the area under the receiver operating characteristic (ROC) curve (AUROC). Paulo et al. (2024) proposed ‘fuzzing’ and ‘detection’ classification tasks to evaluate feature explanations. For ‘fuzzing’ scoring, both positive and negative examples of tokens (i.e., with non-zero and zero activation values, respectively) for a given latent are delimited with special characters, and a language model is prompted to identify which examples have been correctly delimited for the latent given its explanation. For ‘detection’, a language model is asked to identify which examples contain activating tokens for each feature. Bills et al. (2023) originally proposed ‘simulation’ scoring, based on the correlation between predicted and observed activations, but this method is expensive to compute.
@@
Except where noted, we report ‘fuzzing’ scores as a measure of auto-interpretability, because this measure has been demonstrated to correlate with simulation scoring (Paulo et al., 2024). We include similar AUROC curves for the ‘detection’ scoring method in Appendix B. For each trained SAE (i.e., underlying model, variant, and layer), we randomly sampled 100 features to obtain auto-interpretability scores. The implementation is based on Paulo et al. (2024). We use the Meta-Llama-3.1-70B-Instruct-AWQ-INT4 model to generate explanations and make predictions (larger than the 8B models used by Choi et al. (2024) and open-source, unlike Bills et al. 2023).
@@
We found that the auto-interpretability scores were far more similar between the trained and randomized models than with the control (Figures 1 and 2). The similarity between the ROC scores for trained and randomized transformers demonstrates that ‘fuzzing’ auto-interpretability alone, applied to SAE latent explanations, may not meaningfully distinguish between these underlying models.
Evaluation
@@
We considered standard SAE evaluation metrics alongside the auto-interpretability AUROC for Pythia models with between 70M and 6.9B parameters. As above, we broadly found that the randomly initialized and re-randomized models (Figure 2; blue, green, orange lines) were more similar to the trained model (Figure 2; gray lines) than to our control (Figure 2; black lines).
Notably, the cosine similarity between the original activation and the SAE reconstruction and explained variance are often far lower for the random control than the other models, and its reconstruction errors tend to increase across layers while the remaining variants decrease. For the random control, this can perhaps be explained by the fact that a Gaussian is the highest entropy distribution with fixed mean and variance (Jaynes, 2003); we speculate that Gaussian vectors are the ‘least structured’, in some sense, and thus hardest for SAEs to reconstruct. As Gaussian-distributed activations are propagated through successive layers, we would expect the activations to become less Gaussian and perhaps more ‘sparse’, i.e., easier to reconstruct (Section 4).
Interestingly, the randomized variants (blue and orange lines) are more similar to the trained model than the variant at initialization (green line). This is especially evident if we look at the $L^{1}$ norm values in larger models. We speculate that this pattern arises because parameter norms may differ greatly between a trained model and its state at initialization. In contrast, our randomization procedure was specifically designed to preserve parameter norms with respect to the trained model. The scale of parameters at different layers may be important, e.g., to control the growth of activations as they progress through the residual stream (Liu et al., 2020). In the AUROC plots, we find that for all but the control variant, AUROC increases with model size. We speculate that features become more specific as SAE size increases: in smaller SAEs, each latent must explain more of the input, making classification tasks easier for larger SAEs.
Figure 2 (row five) shows the cross-entropy (CE) loss score, or loss recovered, against model layer. This is the increase in the loss when the original model activations are replaced by their SAE reconstructions, divided by the increase when the activations are replaced by zeros (‘ablated’). The results show that the ‘trained’ variant SAEs perform similarly to others from the literature (e.g., Kissane et al., 2024; Rajamanoharan et al., 2024a; Mudide et al., 2024). Importantly, the CE loss score only makes sense for the trained variant: for any of the randomized variants, the loss is very poor, regardless of whether the original or reconstructed activations are used.
Latent explanation complexity
@@
Despite sometimes similar auto-interpretability scores and evaluation metrics, we had expected that SAEs applied to trained vs. randomized transformers would discover qualitatively different features. In particular, we expected SAEs trained on the randomized variants to learn relatively simple features based on characteristics of the input text, but not more complex, abstract features as with trained transformers (Templeton et al., 2024). For qualitative examples, we provide a random sample of features and the corresponding maximally activating dataset examples for each variant of Pythia-6.9b in Appendix J, and more detailed information in Appendix L.
@@
Anecdotally, we have observed that a significant proportion of SAE latents have non-zero activations only on a single token or a small number of distinct tokens within a text dataset (e.g., Lin and Bloom, 2024; Dooms and Wilhelm, 2024). Hence, a simple measure of the complexity of an SAE latent given a set of maximally activating examples is the degree to which the latent activates on a single token ID or multiple distinct IDs. Specifically, we quantify the number of token IDs in terms of the entropy of the observed distribution of latent activations over tokens: the greater the entropy, the more ‘spread out’ the latent activations, and the less token-specific the latent. We take this distribution to be the total latent activation per token across the set of maximally activating examples used to generate explanations for auto-interpretability. We show the relationship between entropy and ‘fuzzing’ AUROC score for individual latents in Appendix H.
@@
We include the entropy of the observed distributions of latent activations over token IDs in the last row of Figure 2. The negative control variant displays a consistently high entropy, which is to be expected given that the embedding for a given token ID is sampled i.i.d. from a Gaussian on each occurrence of the token, i.e., a token does not have a consistent embedding vector (Section 3). For the trained variant, the entropy increases across layers, i.e., the further into the model, the less likely the maximally activating examples for each latent contain activations concentrated on a single token. This is also expected: at later layers, we expect more abstract features that are less similar to token embeddings. Finally, the entropy for randomized models tends to be lower than for either the trained or control variants, indicating that latents are activated specifically at one or a few IDs.
@@
In combination with the preceding results, this suggests that standard SAE quality and auto-interpretability metrics are missing an important aspect of SAE features: their ‘abstractness’. While the token distribution entropy is not a direct measure of ‘abstractness’, it suggests that the randomized variants, viewed in the context of their similar auto-interpretability scores to the trained variant, remain able to learn simple, single-token features. However, unlike the trained variant, the features of the randomized variants do not become more complex as the layer index increases.
4 A toy model of superposition in random networks